The average salary for a certain profession is$ 69,000. Assume that the standard deviation of such salaries is $ 39,500. Consider a random sample of 68 people in this profession and let x over bar x represent the mean salary for the sample.

  

a.

What is mu Subscript x over bar μx​ ?

mu Subscript x over bar μx =________

b.

What is sigma Subscript x over bar σx​ ?

sigma Subscript x over bar σx =__________ ​(Round to two decimal places as​ needed.)

c.

Describe the shape of the sampling distribution of x over bar x.

A. The shape is that of a uniform distribution.

B. The shape is that of a binomial distribution.

C. The shape is that of a poisson distribution.

D. The shape is that of a normal distribution.

d.

Find the​ z-score for the value x over bar x= 59,500.

z=________ ​(Round to two decimal places as​ needed.)

e.

Find Px>59,500.

Px>59,500=_______​(Round to three decimal places as​ needed.)
Answer #1

Solution:- Given that μ = 69000, σ = 39500, n = 68

a. μx = 69000

b. σx = 39500/sqrt(68) = 4790.08

c. option D. The shape is that of a normal distribution

d. Z = (X-μ)/(σ/sqrt(n))

Z = (59500-69000)/(39500/sqrt(68))
  
Z = -1.98

e. P(X > 59500) = P(Z > -1.98) = 0.9761 = 0.976 (rounded)